### v1

@@ Abstract
Figure 6 shows that both the HW and the GAIT models exhibit the expected diffusive scaling $\sigma_{\Delta r}^{2}(t)\sim t$ and Gaussian profiles of $P(\Delta r)$. However, for the parameters used in the results reported above, $(w_{\phi},w_{\nabla\phi},w_{\mathbb{K}\mathbb{L}})=(1,10,0.01)$, labeled as “case 1” in the figure, the value of the diffusivity $D=\sigma_{\Delta r}^{2}/2t$ in the GAIT model, $D_{GAIT,1}=1.28$, tends to be bigger than the diffusivity obtained using the HW model, $D_{HW}=0.933$. This discrepancy can be resolved by using $(w_{\phi},w_{\nabla\phi},w_{\mathbb{K}\mathbb{L}})=(0,10,0.1)$. This case, labeled as “case 2” in the figure, exhibits excellent agreement with $D_{GAIT,2}=0.981$. Most importantly, the GAIT model preserves the same transport properties when extrapolated way beyond the range of training and solution of the HW model.
Figure 6: Comparison of radial displacements, $\Delta r$, of tracers in HW and GAIT models. Left: variance $\sigma_{\Delta r}^{2}(t)$ as function of $t$ in log-log scale. Right: Probability distribution function of displacements at different times.
In the calculations reported in this letter, the numerical integration of the HW model to produce 10000 snapshots of training data took 50 hours on a V100 GPU node of the French Jean-Zay super computer. On the same node, the training of the CVAE took 1 hour, and the training of the DNN less than two minutes. The generation of 10000 snapshots of turbulent states with the GAIT model took about 7.5 min. That is, once trained, the GAIT model can generate turbulent states at a rate 50 hrs/ 7.5 min $\sim$ 400 times faster than the direct numerical integration.
This work has been carried out within the framework of the EUROfusion Consortium, funded by the European Union via the Euratom Research and Training Programme (Grant Agreement No 101052200 ? EUROfusion). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them. This work has received financial support from the AIM4EP project (ANR-21-CE30-0018), funded by the French National Research Agency (ANR), and from the Oak Ridge National Laboratory, managed by UT-Battelle, LLC, for the US Department of Energy under Contract No. DE-AC05-00OR22725. All the simulations and training of neural networks reported here were performed on HPC resources of IDRIS under the allocations 2021-A0100512455, 2022-AD010512455R1 and 2023-A0140514165 made by GENCI.

### v2

@@ Abstract
Case 2, that focuses on the optimization of $\nabla\phi$ (i.e., the ${\bf E}\times{\bf B}$ velocity) and high regularity of the latent space, yields a better description of the Lagrangian transport with a diffusivity $D_{GAIT,2}=1.19$ very close to the ground truth diffusivity $D_{HW}=1.27$, along with a very good agreement on the PDFs of $R$. The smaller value of $w_{\mathbb{K}\mathbb{L}}$ in Case 1, reduces the latent space regularity and leads to a slightly higher value of the diffusivity, $D_{GAIT,1}=1.47$. Most importantly, as proof of the fidelity of the turbulence generation process, the GAIT model preserves the transport properties well beyond ($t=10^{5}$) the numerical integration time of the HW model ($t=8\times 10^{3}$).
Figure 6: Comparison of radial displacement statistics of tracers in the HW and GAIT models . $\langle R^{2}\rangle$ as function of $t$ in log-log scale (left). Probability distribution function of displacements at different times (right).
The direct numerical integration of the HW model for 10,000 snapshots took 50 hours on a V100 GPU node of the French Jean-Zay supercomputer. On the same node, the training of the CVAE took 1 hour, and the training of the RNN five minutes. The generation of 10,000 snapshots of turbulent states with the GAIT model took 7.5 min, about 50 hrs/7.5 min $\sim$ 400 times faster than the direct numerical integration of the HW model.
The HW model requires the integration of the potential, $\phi$, and density, $n$, coupled equations. An advantage of the GAIT model is that it can provide a surrogate model of $\phi$ only, for applications not requiring $n$, like the one discussed here. If needed, a direct generalization of the GAIT model can be used to construct surrogates for $n$ and $\phi$.
This work has been carried out within the framework of the EUROfusion Consortium, funded by the European Union via the Euratom Research and Training Programme (Grant Agreement No 101052200 — EUROfusion). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them. This work has received financial support from the AIM4EP project (ANR-21-CE30-0018), funded by the French National Research Agency (ANR), and from the Oak Ridge National Laboratory, managed by UT-Battelle, LLC, for the US Department of Energy under Contract No. DE-AC05-00OR22725. All the simulations and training of neural networks reported here were performed on HPC resources of IDRIS under the allocations 2021-A0100512455, 2022-AD010512455R1 and 2023-A0140514165 made by GENCI.